A high-speed train travels at 310km/h. How many hours would it take to travel 1500 km at this speed?

Answer by TimothyLamb(4379)

s = d/t
t = d/s
t = 1500/310
t = 4.83870967742 hours
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---